‘When originally purchased, a vehicle had an estimated useful life of 8 years. The vehicle cost $23,000 and it’s estimated salvage value is $1,500. after 4 years of straight-line depreciation, the asset’s total estimated useful life was revised from 8 years to 6 years; there was No change in the estimated salvage value. What is the amount of depreciation expense in 5 years? Show your work

a. $5,375.00

b. $2,687.50

c. $5,543.75

d. $10,750.00

e. $ 2,856.25
Answer #1

Depreciation per year = (Cost - Salvage value) /useful life

= (23,000-1,500)/8 = 2687.50 per year

Book value after 4 years = 23,000 - (2687.50 * 4 years)

= 12,250

Now, useful life is revised to 6 years. But, 4 years have already passed. So, remaining life is 2 years.

Depreciation expense per year for remaining life

= (12,250-1,500)/2

= 5,375